OpenStudy (lgbasallote):

a bacterial culture is known to grow at a rate proportional to the amt. present.after 1 hour, 1000 strands of the bacteria are observed in the culture, and after 4 hours, 3000 strands. find the number of strands of bacteria originally in the culture

OpenStudy (lgbasallote):

at t= 0 x = not given at t = 1 x = 1000 at t = 4 x = 3000

OpenStudy (lgbasallote):

\[\ln(\frac{3000}{1000} = k(4-1)\] \[\ln (3) = 3k\] \[k = 0.3662\]

mathslover (mathslover):

\[\huge{\frac{dA}{dt}=kA}\] where A is the Amount of bacteria and k is a constant. Solve the differential equation and substitute the values to get k and the constant of integration. then put t = 0

mathslover (mathslover):

Source : http://www.mathisfunforum.com/viewtopic.php?id=17940

OpenStudy (lgbasallote):

\[x = x_o e^{0.3672t}\] \[1000 = x_0 e^{0.3662(1)}\] \[1000 = 1.4422x_0\] \[x_0 = 693.36 \implies 694 \]

OpenStudy (anonymous):

you know it triples every 3 hours, so can model as \[1000\times 3^\frac{t}{3}\] where \(t\) is the time starting at hour one to get one hour previous, replace \(t\) by -1

OpenStudy (anonymous):

yikes no need for \(e\) and all that fancy jazz

OpenStudy (anonymous):

no need for differential equations either you are told directly that it triples every three hours, use that immediately

OpenStudy (lgbasallote):

how would you solve it w/o d.e.?

OpenStudy (anonymous):

i just solved it, you can do everything but the last calculation http://www.wolframalpha.com/input/?i=1000*3^%28-1%2F3%29 in your head

OpenStudy (lgbasallote):

wait why \(\large 1000 \times 3^{\frac t3}?\) i didnt quite get your explanation for it sorry

OpenStudy (anonymous):

your are told the growth rate is proportional in the first line in the second line you are told that it triples every three hours

OpenStudy (lgbasallote):

mmhmm

OpenStudy (lgbasallote):

wait..t = 1... x = 1000 t = 3...x = 3000 isnt it per 2 hours?

OpenStudy (anonymous):

so every three hours it triples right? if you start counting at \(t=0\) you have \(p_0\) at \(t=3\) you have \(p_0\times 3\) at \(t=6\) you have \(p_0\times 3^2\) at \(t=9\) you have \(p_0\times 3^3\) and more generally at \(t=t\) you have \(p_0\times 3^{\frac{t}{3}}\)

OpenStudy (anonymous):

you started counting at hour 1, not hour zero

OpenStudy (lgbasallote):

ohh im seeing your logic now

OpenStudy (anonymous):

that is why i had to replace \(t\) by \(-1\) to get back one step

OpenStudy (lgbasallote):

hmm son of a gun..it's right...

OpenStudy (anonymous):

this solving for the rate is a waste of time as far as i am concerned because you are told the rate in the problem triples every three hours

OpenStudy (lgbasallote):

you raise a good point...but it's some out of the box thinking to come up with a relation like that..

OpenStudy (anonymous):

population increases from 100 to 150 every two hours model immediately as \(100\times (1.5)^{\frac{t}{2}}\) no thought required

OpenStudy (anonymous):

no no it is simple thinking keeps me from having to use a different base, when in fact you are told the base the truth is you can use \(e\), \(3\) or \(\pi\) if you like it is just that if you are told the base in the question you do not need to change it to some other base

OpenStudy (lgbasallote):

i see...i'll try that thanks

OpenStudy (lgbasallote):

so im gonna use the difference as base..?

OpenStudy (anonymous):

well your teacher might not like it, but believe me it is fast, works, and is in fact more accurate because you do not truncate as you do when you take the log

OpenStudy (anonymous):

what i mean is this: suppose i am told that the population starts at say 100 and in 2 hours it is 150 that means it increases by 50% every two hours. to increase something by 50% multiply by \(1.5\) which is of course the same as computing \(\frac{150}{100}\) so now i have my base as \(1.5\) and since it increases by 50% every 2 hours i write \[100\times (1.5)^{\frac{t}{2}}\] works the same for half life

OpenStudy (lgbasallote):

ohhh i think im beginning to get the idea

OpenStudy (anonymous):

whereas the other method requires me to solve \[100e^{2k}=150\] for \(k\) and then go back to write my formula

OpenStudy (lgbasallote):

hmm i agree that diff eq is long..

OpenStudy (anonymous):

my car ( a nice old satellite) depreciates from 1200 to 800 in 3 years i model immediately as \(1200\times (\frac{2}{3})^{\frac{t}{3}}\)